OpenStudy (anonymous):

i need help with my math 8th grade work please help What is the solution to the system of equations represented by the two equations? Y=2/3X Y=2/3X+4 please help

OpenStudy (anonymous):

are the equations y=2/(3x) and y=2/(3x+4)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

The system don't have any solution.

OpenStudy (anonymous):

@boygame20